OpenStudy (anonymous):

if tan A= -2/3 and tan B = 9/4 find tan(A-B)

jimthompson5910 (jim_thompson5910):

tan(A-B) = ( tan(A) - tan(B) )/( 1 + tan(A)*tan(B) ) tan(A-B) = ( -2/3 - 9/4 )/( 1 + (-2/3)(9/4) ) tan(A-B) = ( -2/3 - 9/4 )/( 1 - 18/12 ) tan(A-B) = ( -2/3 - 9/4 )/( 1 - 3/2 ) tan(A-B) = ( -35/12 )/( -1/2 ) tan(A-B) = ( -35/12 )*( -2/1 ) tan(A-B) = 35/6